### v1

# Street Gaussians without 3D Object Tracker

Abstract
Realistic scene reconstruction in driving scenarios poses significant challenges due to fast-moving objects. Most existing methods rely on labor-intensive manual labeling of object poses to reconstruct dynamic objects in canonical space and move them based on these poses during rendering. While some approaches attempt to use 3D object trackers to replace manual annotations, the limited generalization of 3D trackers – caused by the scarcity of large-scale 3D datasets – results in inferior reconstructions in real-world settings. In contrast, 2D foundation models demonstrate strong generalization capabilities. To eliminate the reliance on 3D trackers and enhance robustness across diverse environments, we propose a stable object tracking module by leveraging associations from 2D deep trackers within a 3D object fusion strategy. We address inevitable tracking errors by further introducing a motion learning strategy in an implicit feature space that autonomously corrects trajectory errors and recovers missed detections. Experimental results on Waymo-NOTR datasets show we achieve state-of-the-art performance. Our code will be made publicly available.
Figure 1: Comparison of 3D tracker-based Street Gaussians [76] (left) and our approach (right). Existing methods heavily rely on object poses, but 3D trackers struggle with limited generalization, leading to flaws in novel view synthesis. In contrast, 2D foundation models show better generalization. Our approach leverages a 2D foundation model [72] for object tracking and learns point motion within an implicit feature space to autonomously correct tracking errors, improving robustness across diverse scenes.
1 Introduction
Modeling dynamic 3D street scenes underpins modern autonomous driving by enabling realistic, controllable simulations for tasks such as perception [46, 11, 81, 38], prediction [30, 42, 22], and motion planning [13, 14, 16]. With the rise of end-to-end autonomous systems that require real-time sensor feedback [26, 28, 31], high-quality scene reconstructions have become essential for closed-loop evaluations [83, 41], particularly to simulate critical corner cases safely and cost-effectively.
Despite extensive efforts in achieving photo-realistic reconstruction of small-scale scenes, the unique challenges posed by the large-scale and highly dynamic nature of driving scenarios complicate effective 3D scene modeling. To address these challenges, most existing methods [34, 73, 88, 55] rely on ground truth vehicle poses to differentiate between static background and moving vehicles. Typically, vehicles are reconstructed in canonical space and subsequently positioned based on known poses during rendering. However, collecting ground-truth poses is labor-intensive, limiting the applicability of these methods to scenes beyond the existing datasets.
To eliminate reliance of the systems on ground truth object poses, Street Gaussians [76] instead uses poses generated by 3D object trackers [70, 71]. However, by defining vehicle motion solely through these tracking poses, rendering quality becomes highly dependent on pose accuracy. While Street Gaussians optimizes object poses during training, it struggles with detection failures and large pose errors, as shown in Fig. 1. Unfortunately, 3D trackers often struggle to generalize effectively across different scenarios, primarily due to the scarcity of open-source large-scale 3D datasets [86, 67]. Collecting ground-truth poses for 3D tracking is time-consuming and costly, resulting in relatively few open-source datasets for autonomous driving. For example, even widely-used datasets such as Waymo [65], Nuscenes [6], KITTI [21, 44], Pandaset [74] collectively contain fewer than 4,000 annotated scenes, limiting the diversity and scale needed for robust 3D tracking. In contrast, 2D data is far easier to collect and annotate, leading to an abundance of 2D datasets. The BDD100K dataset [82] alone includes annotations for 100,000 driving scenes, and additional large-scale 2D perception datasets, such as COCO [45], LVIS [25], and OpenImages [37], bring the total number of annotated scenes into the millions. This wealth of 2D data has enabled 2D foundation models to achieve strong generalization across a variety of tasks, including embedding extraction [54, 57], detection [52, 8], segmentation [36, 12], and tracking [72]. Nonetheless, how to use these accurate and robust 2D trackers to effectively model dynamic objects in 3D is still an open challenge for street scene reconstruction.
In this paper, a novel architecture is designed to achieve high-fidelity novel view synthesis performance in street scenarios for autonomous driving applications. Unlike previous approaches such as Driving Gaussian [88] and Street Gaussian [76] that rely on ground truth vehicle poses or poses predicted by 3D trackers, the first contribution of this work is a stable object tracking module. This module leverages associations from 2D deep trackers within a 3D object fusion framework to enhance robustness and accuracy. Specifically, we integrate 2D tracking outputs with LiDAR data to trace the trajectory of each vehicle in 3D. We then incrementally reconstruct the point cloud of each vehicle frame-by-frame in the canonical space, estimating its pose by aligning successive frames with this canonical model. This approach eliminates the reliance on 3D trackers and enhances robustness across diverse environments.
Although 2D tracking models demonstrate stronger generalization ability, tracking errors still can exist, especially under adverse conditions such as severe occlusions or distant objects. Since moving the object points solely based on the tracked object pose directly exposes any tracking errors, we aim to go beyond a straightforward reliance on tracked trajectories. Instead, we propose to learn point motion from the predicted trajectory, equipping the model to autonomously identify and correct tracking inaccuracies, recover missed detections, and infer motion in new time steps. An implicit representation is essential for this purpose: 1) It enables the model to refine trajectories without being constrained by bounding box tracks, facilitating smoother and more continuous motion that can adaptively respond to changes. 2) An implicit feature space offers versatility to move each point of an object in a different way, recognizing that vehicles are not strictly rigid objects — for example, doors can open or close. This approach makes it possible to capture subtle, dynamic changes within objects, ultimately producing more accurate reconstructions and enhancing the robustness of novel view synthesis in challenging scenarios. To this end, we leverage HexPlane representation [7] following 4DGS [69] and apply direct supervision on learned point motion using the predicted trajectory. However, applying 4DGS alone without the object tracking module and explicit motion supervision yields suboptimal results due to two main challenges: 1) Reconstructing fast-moving objects without tracking is difficult since aggregating points from different time frames results in a fragmented trail rather than a smooth and coherent representation. 2) Modeling rapid vehicle motion without explicit supervision proves challenging. For example, the projection of a Gaussian may fall outside the actual object area when the initial position is distant from its current position, impeding optimization since gradients cannot propagate from the object area to the Gaussian. These issues hinder precise motion modeling, as seen in S3Gaussian [29], which uses 4DGS without explicit supervision and thus struggles with moving cars, as demonstrated in our experiments.
Our main contributions are summarized as follows:
  • We eliminate the need of 3D tracker for street scene reconstruction by introducing a robust object tracking module which leverages 2D foundation model, achieving superior generalization ability across diverse scenarios.
  • We introduce a motion learning framework to learn from the predicted trajectory in an implicit feature space, enabling it to automatically correct pose errors and infer motion for novel time steps.
  • We achieve state-of-the-art performance on Waymo-NOTR dataset without relying on ground truth annotations.
2 Related Works
2.1 3D Gaussian Splatting for Dynamic Scene
3D Gaussian Splatting (3DGS) [33] has advanced scene reconstruction by enabling high-quality, real-time rendering with 3D Gaussians and efficient splat-based rasterization, reducing computation and parameters compared to NeRF-based methods [51, 1, 2, 3, 53] and other representations [15, 23, 64, 24, 27, 40]. While originally designed for static scenes, 3DGS has been adapted for dynamic scenes [47, 79, 80, 69, 63, 43]. Dynamic3DGS [47] directly stores information for each 3D Gaussian at every timestamp, and Yang et al. [79] approximate the spatiotemporal 4D volume by optimizing 4D Gaussian primitives. Deformable-3DGS [80] and GauFRe [43] employ deformation fields to model motions, while 4DGS [69] introduces HexPlane representation [7] to store spatial-temporal features efficiently. Gaussian Marbles [63] leverages isotropic “marbles” and a divide-and-conquer trajectory learning algorithm. Despite these advances, challenges persist in high-speed driving scenarios with rapid object motion. To tackle this challenge, we leverage a robust object tracking module to guide the motion learning process.
2.2 Street Scene Reconstruction
Existing autonomous driving simulation engines [19, 60, 39] face high manual effort in creating virtual environments and a lack of realism in the generated data. The creation of high-fidelity simulations from driving logs is therefore essential for advancing closed-loop training and testing. Recent works [78, 75, 55, 66, 61, 73, 77] have continuously made improvements to NeRF [51] to model street scenes dynamically. Despite the progress, NeRF-based methods remain computationally expensive and require densely overlapping views. Building on the effective 3DGS approach [33] for scene reconstruction, several 3DGS-based methods have emerged. PVG [10] model dynamic scenarios by using periodic vibration-based temporal dynamics. Driving Gaussian [88] introduces incremental static Gaussians and composite dynamic Gaussian graph. Street Gaussians [76] equips Gaussians with semantic logits, and optimize dynamic parts using tracked poses from 3D tracker. AutoSplat [34] enforces geometric constraints in road and sky regions for multi-view consistency. Among these approaches, most of them requires ground truth object pose [10, 88, 34] or the 3D tracker [76]. However, manual annotation is laborious and the 3D tracker is lack in generalization ability, which limits their applications in diverse scenarios. In contrast, S3Gaussian [29] models dynamics in a self-supervised manner using dynamic Gaussians from 4DGS [69], despite struggling with dynamic object modeling due to the lack of explicit motion supervision. We address this challenge by introducing a robust object tracking strategy based on a 2D foundation model [72] and apply motion supervision from the predicted object trajectory.
2.3 2D and 3D trackers
Multiple Object Tracking (MOT) aims at locating multiple objects in each frame, and establishing correspondences between them across frames within input videos [4, 85, 48, 32, 5]. The reliance of most MOT methods on object detection [4, 85, 48, 56, 18, 87] makes detection accuracy crucial for MOT performance. Recent advances in 3D object detection are promising [68, 50, 81, 11, 46, 38], but these models often generalize poorly due to limited and biased datasets [62, 84, 20]. Existing open-source autonomous driving datasets lack scale and show regional or environmental biases, such as vehicle density and weather conditions [65, 6, 21, 49]. The main reason is that building large-scale, well-annotated multimodal datasets is costly [6, 49, 58, 44]. In contrast, 2D image data is easier to capture and annotate, enabling extensive dataset collections [9, 82, 17, 37]. Visual foundation models trained on large-scale 2D data have shown strong generalization [72, 12, 57, 36, 54]. We thus opt to use 2D object tracker [72] to locate the dynamic objects.
Figure 2: Overview of our method. To overcome the limited generalization of 3D object trackers, we introduce a robust object tracking module based on a 2D object tracker [72]. We integrate 2D tracking with LiDAR data to segment the object’s point cloud and incrementally reconstruct the object model in canonical space frame-by-frame. The canonical model is used to estimate object pose $\mathbf{T}_{t}$ and serves as the initialization for Gaussian optimization. We model dynamic objects with isotropic Gaussian marble [63] to simplify optimization. Our approach learns point motion within the HexPlane [7] feature space, enabling tracking error correction and missed detection recovery. Decoders $D_{\mathcal{X}},D_{\mathcal{C}}$ predict point motion $\Delta\mathcal{X}_{t}$ and color change $\Delta\mathcal{C}_{t}$ based on the HexPlane features $f(x,y,z,t)$, with the predicted trajectory providing supervision. Finally, novel view synthesis is performed by splatting the deformed Gaussian onto the image plane.
3 Our Method
3.1 Prerequisites: 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [33] represents a 3D scene explicitly as a collection of 3D Gaussian primitives. Each Guassian is defined by its center position $\mathcal{X}\in\mathbb{R}^{3}$ and covariance $\Sigma\in\mathbb{R}^{3\times 3}$. To enforce semi-positive definiteness, the covariance matrix is further factorized into a scaling vector $S\in\mathbb{R}^{3}$ and a rotation matrix $R\in SO(3)$ by:
$\Sigma=RSS^{\top}R^{\top}$ (1)
Additionally, each Gaussian is described with its opacity $o\in\mathbb{R}$ and view-dependent color defined by spherical harmonic (SH) coefficient $\mathcal{C}\in\mathbb{R}^{k}$, where $k$ represents numbers of SH functions.
During rendering of novel views, differential splatting is applied to 3D Gaussians within the camera planes. The blending of $N$ ordered points that overlap a pixel is given by the formula:
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
where $\alpha_{i}$ and $c_{i}$ represents the opacity and color of the $i$th splatted Gaussian, which is computed from per-point opacity and SH coefficients (see [33] for details).
3.2 Overview
Our method takes as input multi-view images $\mathbf{I}_{t}^{(j)}$ from multiple cameras placed around the vehicle, each indexed by time step $t$ and camera index $j$, along with intrinsic $\mathbf{K}^{(j)}$ and extrinsic $\mathbf{E}_{t}^{(j)}$ matrices for each view. Additionally, a top-mounted LiDAR provides 3D point clouds $\mathbf{L}_{t}$ for each frame. Using this multi-sensor data, our approach reconstructs the 3D scene and synthesizes novel views from any desired camera pose and time frame, without reliance on 3D object trackers or ground truth object trajectories.
We present an overview of our method in Fig. 2. Visual images and point clouds are obtained from the sensor setup. For images, static and dynamic components are analyzed first. We employ Mask2Former [12] to segment the scene into static and dynamic parts. Specifically, the dynamic part includes humans (pedestrians, cyclists, etc.) and vehicles (cars, trucks, etc.), and all the other objects are categorized as static. We model the static part as in 3DGS [33] and focus on modeling dynamic objects in the following sections. To enhance robustness, we propose an object tracking strategy by leveraging a robust 2D tracker [72] and the LiDAR point cloud, providing accurate initial point clouds and object trajectories for subsequent Gaussian optimization (cf. Sec 3.3). We leverage an implicit representation HexPlane [7] following 4DGS [69] to learn a continuous and smooth per-point motion based on the object trajectory (cf. Sec 3.4). The optimization objective is described in Sec 3.5.
3.3 Tracking Cars in 3D with Robust 2D Tracker
To improve the robustness of novel view synthesis for diverse scenarios, the first challenge is how to track and build vehicles in 3D based on the 2D tracking results.
We employ GLEE [72] to track all vehicles within the 2D image plane, obtaining 2D object trajectories in each camera view. Each trajectory includes the 2D segmentation masks of the object over a time period. To lift 2D trajectories into 3D, we first re-project the LiDAR point clouds onto the image plane and assign points to objects based on their presence within the corresponding 2D segmentation masks $\mathcal{M}$. This yields a segmented object point clouds for each camera view via the following function:
$\mathcal{O}_{t}^{(j)}=\{\mathbf{P}|\Pi^{-1}(\mathbf{P},\mathbf{K}^{(j)},\mathbf{E}^{(j)}_{t})\in\mathcal{M},\mathbf{P}\in\mathbf{L}_{t}\},$ (3)
where $j$ is the camera index, $t$ is the time index, $\mathbf{P}$ is a LiDAR point and $\Pi(\cdot)$ is the re-projection function. Given the slight misalignment between the LiDAR point cloud and RGB image, points near the edges of segmentation masks often fall outside the object boundaries. To address this issue, we perform outlier removal to improve stability in subsequent reconstruction based on the point distance to the point cloud center. We then associate the same objects across different camera views. Two point clouds in different viewpoints are considered belonging to the same object if they share more than 50 points in one time step. This association process results in a set of associated partial object point cloud in different time steps.
After obtaining the 3D partial objects in different time steps, the next step is to associate these partial point cloud into a unified and complete model. We initialize the 3D reconstruction of each object from the first frame which it appears. For simplicity, we assume an object is visible between frames 00 and $T$ and denote the object point cloud at time $t$ as $\mathcal{O}_{t}$. Our goal is to obtain a temporally consistent reconstruction $\mathcal{O}$ and the object pose in each time frame. This reconstruction is done frame-by-frame incrementally.
Starting with the initial frame, we add $\mathcal{O}_{0}$ to $\mathcal{O}$. For each subsequent frame, we apply Iterative Closest Point (ICP) for the alignment of $\mathcal{O}_{t}$ to $\mathcal{O}$ by extracting the relative pose $\mathbf{T}_{t}$. The overlap between $\mathcal{O}_{t}$ and $\mathcal{O}$ is given by:
$\tau_{overlap}=\mathcal{Q}\cap\mathbf{T}_{t}^{-1}\mathcal{O}_{t},\quad\mathbf{T}_{t}=\operatorname{ICP}(\mathcal{O}_{t},\mathcal{O}).$ (4)
If ICP reveals an overlap $\tau_{overlap}$ of more than 30% with $\mathcal{O}$, we record the object pose $\mathbf{T}_{t}$ and update $\mathcal{O}$ by integrating $\mathbf{T}_{t}^{-1}\mathcal{O}_{t}$. If $\tau_{overlap}$ is between 10% and 30%, we only record the object pose $P_{t}$ without updating $\mathcal{O}$. If $\tau_{overlap}$ is lower than 10%, we assume a 2D tracking failure for that frame and discard it to filter out tracking errors. This process is repeated for each frame using the last recorded pose as the initial guess for ICP to accelerate convergence. Through this approach, we reconstruct all objects and retrieve their poses in each frame, providing a robust initialization for subsequent 3D Gaussian optimization.
3.4 Learning Point Motion
Task Scene Reconstruction Novel View Synthesis
Method Extra Input PSNR↑ SSIM↑ LPIPS↓ DPSNR↑ DSSIM↑ PSNR↑ SSIM↑ LPIPS↓ DPSNR↑ DSSIM↑
3DGS [33] N/A 25.77 0.833 0.160 20.26 0.604 24.46 0.802 0.170 18.12 0.521
EmerNeRF [77] N/A 28.16 0.806 0.228 24.32 0.682 25.14 0.747 0.313 23.49 0.660
S3Gaussian [29] N/A 31.35 0.911 0.106 26.02 0.783 27.44 0.857 0.137 22.92 0.680
MARS [73] GT Pose 28.24 0.866 0.252 23.37 0.701 26.61 0.796 0.305 22.21 0.697
StreetGS [76] 3D tracker 29.17 0.873 0.138 27.78 0.818 26.98 0.838 0.149 24.62 0.742
Ours 2D tracker 32.56 0.936 0.059 28.51 0.868 28.85 0.867 0.088 25.58 0.779
Table 1: Comparison with state-of-the-art methods on NOTR dataset. StreetGS represents Street Gaussian [76]. The best and the second best results are denoted by pink and blue.
A common approach to model object motion in prior works is directly transforming the object points using the given object pose. However, despite the strong generalization capabilities of 2D trackers, detection failures are inevitable in challenging scenarios, such as when objects are heavily occluded or located at great distances. Relying solely on object poses derived from these erroneous or missing detections can easily lead to rendering failures. Additionally, treating cars as rigid objects does not adequately address corner cases, such as when a car door is open or closed. Furthermore, these methods lack the ability to infer an object motion at arbitrary time stamps.
To address these limitations, we seek for a motion modeling approach to enhance robustness and flexibility. Instead of explicitly using object pose as rigid transformation, we learn the per-point motion in a pre-defined feature space. In this feature space, the object motion can be optimized through both explicit guidance and photometric loss. Moreover, the object motion can be interpolated through time and space in this feature space to compensate for missing detections. To this end, we leverage HexPlane representation [7] as in 4DGS [69] to efficiently capture spatial and temporal information by decomposing the 4D feature voxels into six learnable feature planes. This representation satisfies all our requirements and is also memory efficient.
We use three planes: $P_{xy}$, $P_{yz}$, and $P_{xz}$ for the spatial dimensions, and another three planes $P_{xt}$, $P_{yt}$, and $P_{zt}$ for the spatial-temporal features. Additionally, the Hexplane includes multiple resolution levels which is formulated as:
$\begin{split}\{P_{ij}^{\rho}\in\mathbb{R}^{d\times\rho{r_{i}}\times\rho{r_{j}}}|(i,j)\in\mathcal{P},\,\rho\in\{1,2\}\},\end{split}$ (5)
where $\mathcal{P}=\{(x,y),(x,z),(y,z),(x,t),(y,t),(z,t)\}$, $d$ is the feature dimension, $\rho$ denotes the upsampling scale, and $r$ is the base resolution.
Given the center position of a Gaussian $(x,y,z)$ and the time step $t$, features in all six planes are queried and combined via a small MLP $\phi_{d}$ as follows:
$f(x,y,z,t)=\phi_{d}(\bigcup_{\rho}\prod_{(i,j)\in\mathcal{P}}\pi(P_{ij}^{\rho},\psi_{ij}^{\rho}(x,y,z,t))),$ (6)
where $\psi_{ij}^{\rho}(x,y,z,t)$ projects the 4D coordinate onto the respective plane, and $\pi$ performs bilinear interpolation on the voxel features at each point. The features of all planes are combined using the Hadamard product.
Finally, two MLP decoders $D_{\mathcal{X}},D_{c}$ are utilized to predict the point motion $\Delta\mathcal{X}_{t}$ and color change $\Delta\mathcal{C}_{t}$ as $\Delta\mathcal{X}_{t}=D_{\mathcal{X}}(f(x,y,z,t))$ and $\Delta\mathcal{C}_{t}=D_{\mathcal{C}}(f(x,y,z,t))$, respectively. The deformed 4D Gaussians are formulated as: $\mathcal{G}_{t}=\{\mathcal{X}+\Delta\mathcal{X}_{t},\mathcal{C}+\Delta\mathcal{C}_{t},S,R,o\}$.
To allow the HexPlane feature to learn point motion from the predicted pose, we define the motion loss as:
$\mathcal{L}_{\text{motion}}=\operatorname{avg}_{\mathcal{X}\in\mathcal{O}}|\Delta\mathcal{X}_{t}-(\mathbf{T}_{t}\mathcal{X}-\mathcal{X})|,$ (7)
where $\mathcal{X}$ is the center position of a Gaussian in object $\mathcal{O}$ and $\operatorname{avg}$ is the average operator. This loss encourages the predicted point motion $\Delta\mathcal{X}_{t}$ to align with the predicted pose. We apply this loss only for the first 40% iterations to provide a strong initial motion prior. The loss is subsequently removed, allowing the network to adjust and potentially compensate for pose errors and detection failures.
Following [63], we adopt isotropic Gaussian marbles for dynamic points to reduce degrees of freedom and simplify optimization. In this setup, the rotation of each Gaussian is represented by the identity matrix with identical scales across all three dimensions. The spherical harmonics coefficients are also limited to three dimensions, providing view-consistent color. This approach ensures that all deformations are captured purely by point motion and color changes, enforcing strong constraints and enhancing robustness in novel view synthesis.
Ours vs. S3Gaussian. Although S3Gaussian [29] also uses Hexplane representation [7] and the deformation network from 4DGS [69], our design serves a different purpose: S3Gaussian learns point motion without extra supervision, while we use 4DGS to refine motion and address detection failures. Consequently, S3Gaussian becomes susceptible to unsatisfactory results with rapid car motion due to the lack of explicit guidance. In contrast, our approach enables high-quality reconstruction of dynamic objects.
3.5 Optimization Objective
Besides the motion loss introduced above, the loss function comprises five components to collectively optimize the scene representation and the point motion. The primary loss $\mathcal{L}_{\text{rgb}}$ is an L1 loss measuring the photometric difference between rendered and ground truth images while $\mathcal{L}_{\text{ssim}}$ evaluates their structural similarity. The L1 loss between the rendered depth map and the depth estimated from the LiDAR point cloud $\mathcal{L}_{\text{depth}}$ is used to supervise the Gaussian positions. Following K-Planes [59], a grid-based total variation loss $\mathcal{L}_{\text{tv}}$ is introduced. Recognizing that the color of most dynamic points in the scene are unchanged, a L1 regularization loss $\mathcal{L}_{\text{color-reg}}$ is applied to the deformation network by minimizing $\Delta\mathcal{C}$.
The total loss function is thus defined as:
$\begin{split}\mathcal{L}=&\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}}+\lambda_{\text{tv}}\mathcal{L}_{\text{tv}}\\ &+\lambda_{\text{color-reg}}\mathcal{L}_{\text{color-reg}}+\lambda_{\text{motion}}\mathcal{L}_{\text{motion}},\end{split}$ (8)
where the weights are assigned as follows: $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{depth}}=1.0$, $\lambda_{\text{ssim}}=0.1$, $\lambda_{\text{tv}}=0.1$, $\lambda_{\text{color-reg}}=0.01$ and $\lambda_{\text{motion}}=1$.
The optimization process is divided into two stages. In the first stage, we render using only static Gaussians and supervise only the static regions of the images over 20,000 iterations. This phase focuses on reconstructing the static background and stabilizes the training process. In the second stage, we train all Gaussians on the whole images for 50,000 iterations.
4 Experiments
Figure 3: Qualitative comparison of novel view synthesis on NOTR dataset. Best viewed with zoom.
4.1 Experimental Setup
Datasets. The NeRF On-The-Road (NOTR) dataset, introduced by EmerNeRF [77], is a subset of the Waymo Open dataset [65] and includes various challenging driving scenarios, such as high-speed, exposure mismatch, dawn/dusk, gloomy, rainy, and night scenes. In contrast, many public LiDAR datasets such as nuScenes [6] and KITTI [44] mainly features simpler scenes with limited dynamic content. We leverage NOTR dynamic32 subsets, comprising 32 dynamic scenes to ensure a diverse benchmark for evaluating dynamic reconstruction.
Implementation Details. We use the Adam optimizer [35] with the same learning rate schedule as in 3DGS [33]. For long-sequence scene reconstruction, we follow S3Gaussian [29] and segment the scene into multiple clips. The multi-resolution HexPlane encoder has a base resolution of 64, upsampled by factors of 2 and 4 as in [69], and other hyperparameters match those in 3DGS [33]. As LiDAR points lack data for the sky region, we add a plane of points above the scene’s maximum height to represent it following [34]. For the Waymo-NOTR dataset, we strictly follow EmerNeRF’s experimental setup [77]. For camera images, we use three frontal cameras and resize them to a resolution of $960\times 640$ for both training and evaluation. For scene reconstruction, all samples within a log are utilized for training, probing the upper bound of each method’s performance. For novel view synthesis, we exclude every 10th time step, creating a 10% set of novel temporal views for evaluation. All experiments are conducted on a single NVIDIA RTX 3090 GPU, with training taking approximately 2 hours per video clip. The inference speed is around 80 FPS in resolution $960\times 640$.
Baseline Methods. We evaluate our approach against state-of-the-art methods, covering both NeRF-based and Gaussian-based 3D models. The NeRF-based methods include MARS [73] and EmerNeRF [77]. Gaussian-based methods include 3DGS [33], Street Gaussians [76] and S3Gaussian [29]. To ensure fair comparisons, we apply LiDAR point cloud initialization and add depth regularization to 3DGS. We borrow results of MARS [73] and EmerNeRF [77] from S3Gaussian [29]. Since Street Gaussians requires 3D tracking results as input, we evaluate it using detection results from VoxelNext [11] and associate the detections with SimpleTrack [56]. As NOTR is part of the training set for the perception task in the Waymo dataset, we use VoxelNext pretrained on the nuScenes dataset [6]. This setup also demonstrates the generalization capabilities of methods relying on 3D trackers.
Metrics. We leverage peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) and learned perceptual image patch similarity (LPIPS) to evaluate the rendering quality. Additionally, following EmerNeRF [77], we apply DPSNR and DSSIM metrics to dynamic objects by projecting their ground truth 3D bounding boxes onto the 2D image plane and calculating pixel loss within these regions.
4.2 Comparisons with the State-of-the-art
On the NOTR dataset, our method outperforms all competitors across every metric, as shown in Table 1. Our approach sets a new state-of-the-art in both the scene reconstruction and novel view synthesis tasks. Specifically, our method outperforms S3Gaussian [29] by 2.66 dB in DPSNR and 0.099 in DSSIM for novel view synthesis. It also surpasses Street Gaussians [76] by 1.87 dB in PSNR, mainly due to the limited generalization of 3D trackers used in Street Gaussians. Despite optimizing object poses, Street Gaussians struggles with large pose errors and detection failures, leading to inferior performance. Additionally, we outperform NeRF-based methods like EmerNeRF and MARS.
We present qualitative comparison with S3Gaussian [29] and Street Gaussians [76] in Fig. 3. Results from S3Gaussian show that using 4DGS without explicit motion guidance results in weaker performance when handling moving vehicles (e.g. row 2, 3, 5). Street Gaussians suffers from tracking errors of 3D trackers, leading to erroneous reconstructions (e.g. row 1, 3) or entirely missed objects (e.g. rows 2, 4, and 5). In contrast, our approach performs robustly across diverse scenarios owing to our tracking strategy leveraging 2D foundation model and robust motion learning module.
4.3 Initial Reconstruction Quality Evaluation
Figure 4: Histogram of reconstruction errors of our method and the 3D tracker-based reconstruction. All results with errors greater than 0.6 $m$ are placed in the rightmost bin.
To further evaluate the reconstruction quality of our object tracking module introduced in Sec. 3.3, we conduct an experiment on NOTR dataset. We use the ground truth 3D bounding box to crop the LiDAR point cloud within the box and transform it into the canonical space as the reference reconstruction. Only moving objects that are visible in at least two frames and contain more than 1,000 points in total are preserved. To associate ground truth objects with our predictions, we match each ground truth object with the nearest predicted object across all frames. Since our reconstruction is not in the same canonical space as the reference, we do the alignment before evaluation. For comparison, we repeat the same reconstruction process with 3D bounding boxes predicted by VoxelNext [11] as our baseline. We use Chamfer Distance to evaluate the reconstruction error.
The results are displayed as a histogram in Fig. 4. Note that all results with an error larger than 0.6 meters are placed in the rightmost bin. It can be seen that our method achieves lower reconstruction errors compared to the 3D tracker-based reconstruction. Specifically, 28.3% of the data points in the 3D tracker-based reconstruction have errors greater than 0.5 meters, while all data points in our method have errors below 0.5 meters, highlighting the robustness of our method.
4.4 Ablation Studies
No. Method PSNR↑ SSIM↑ DPSNR↑ DSSIM↑
A w/o track 25.25 0.761 21.71 0.662
B w/o $\mathcal{L}_{\text{motion}}$ 28.39 0.843 24.54 0.719
C w/o iso. 28.58 0.847 25.35 0.765
D 3D tracker 27.55 0.829 24.85 0.747
E GT pose 28.98 0.873 25.71 0.787
F Ours 28.85 0.867 25.58 0.779
Table 2: Ablation study of Novel View Synthesis on NOTR dataset.
We conduct an ablation study on the NOTR dataset and the results are presented in Tab. 2.
Effect of object tracking module. In Tab. 2 (A), we model dynamic points using 4DGS [69] without performing any form of tracking, and observe significant drops across all metrics. In Tab. 2 (B), we use the reconstruction of the tracking module for initialization but remove the vehicle motion loss introduced in Eq. 7, resulting in a 1.17 dB drop in DPSNR. These experiments highlight that the sole use of 4DGS without the support of our tracking module, does not provide sufficient accuracy for effective motion modeling in autonomous driving scenarios.
Effect of isotropic Gaussian marbles. In Tab. 2(C), we substitute isotropic Gaussian marbles with the original anisotropic Gaussian ellipsoids for dynamic points, which leads to a decrease in DPSNR and DDSIM.
3D tracker / GT pose vs. Our object tracking module. In Tab. 2 (D), we use the 3D object tracker [11, 56] to reconstruct dynamic objects and supervise the motion learning, yielding a 1.3dB decrease in PSNR. This suggests that tracking errors can impact novel view synthesis performance despite our motion learning design, highlighting the necessity of substituting 3D tracker with more generalizable 2D tracker. In Tab. 2 (E), we conduct an experiment using the ground truth object trajectory to benchmark our 2D tracking approach. The relatively small performance gap indicates that our 2D tracking closely approximates the ground truth in our motion learning procedure, supporting robust performance across varied scenarios.
Effect of motion learning module. Compared to Street Gaussians which uses the same tracking results in Tab. 1, the PSNR in Tab. 2 (D) is 0.57 dB higher, indicating the robustness of our motion tracking module to tracking errors.
5 Conclusion
In this paper, we introduce a novel framework for robust dynamic 3D street scene reconstruction that eliminates the reliance on 3D object trackers. Addressing the generalization limitations of 3D trackers, we propose a robust object tracking strategy based on a 2D foundation model. Our framework also features a motion learning module within an implicit feature space to handle inevitable tracking errors by autonomously refining pose inaccuracies and recovering missed detections. Experiments on the Waymo-NOTR dataset demonstrate its adaptability and superior performance, underscoring its effectiveness in producing high-quality 3D reconstructions across diverse scenarios.